OpenStudy (anonymous):

Question: Fundamental Theorem of Calculus: We already have a definition of [a,b], it it makes no sense if the interval is opened (has to be closed). How would you go about saying that it has to be on a closed interval?

OpenStudy (amistre64):

i dont think it has to be restricted to a closed interval; hence, improper integrals

OpenStudy (amistre64):

\[\int_{0}^{\infty}tan^{-1}(x)dx\]

OpenStudy (anonymous):

so it doesn't have to be on a closed interval? hmmm... makes sense

OpenStudy (amistre64):

it is useful on a closed interval thats fer sure; but "need to be"? nah

OpenStudy (amistre64):

i think I got my example wrong, but the premise is still the same ;)

OpenStudy (anonymous):

do you have a better example? :)

OpenStudy (amistre64):

sure \[\int_{1}^{\infty}\frac{1}{x^2}dx=1\] http://www.wolframalpha.com/input/?i=integrate+%281+to+infinity%29+1%2Fx%5E2

OpenStudy (anonymous):

Thank YOU!!

OpenStudy (amistre64):

youre welcome